OpenStudy (anonymous):

Sketch the region enclosed by the lines x = 0, x = 6, y = 2, and y = 6. Find the area of the region.

OpenStudy (anonymous):

one side is 6-0=6 and the other is 6-2=4 and 6*4=24

OpenStudy (anonymous):

Thank you @goobersmith :)